OpenStudy (anonymous):

12 = -4(-6x-3) Solve the equation. Would I distribute the -4 to the stuff inside the parentheses?

OpenStudy (anonymous):

Yes, you would so -4 x -6x = ? -4 x-3 = ?

OpenStudy (anonymous):

How would that work with the -6x?

OpenStudy (anonymous):

12 = -4(-6x-3) 12= 24x+12

OpenStudy (anonymous):

a Negative multiplied by a negative = a positive

OpenStudy (anonymous):

x = 0?

OpenStudy (anonymous):

Yes, so there is no solution.

OpenStudy (anonymous):

12=24x+12 subtract 12 from both sides 0=24x divide both sides by 24 0=x

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

No Problem :)